OpenStudy (anonymous):

-6x+7y=20 2x-3y=4 what is the y intersection?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

The y intercept you mean? Set up the equations in y=mx+b format. b is the y intercept. 1. 7y= 6x+20, y = 6/7x + 20/7, 20/7 is the y intercept. 2. 2x-3y=4, -3y=-2x+4, y=2/3x-4/3, 4/3 is the y intercept.